OpenStudy (anonymous):

Evaluate the difference quotient for the given function. f(x) = 1/x, [f(x) - f(a)]/(x-a) I don't understand how you start this off at all. I know that f(a) = 1/a but I don't understand how you get that, either.

jimthompson5910 (jim_thompson5910):

\[\Large \frac{f(x) - f(a)}{x-a}\] \[\Large \frac{\frac{1}{x} - \frac{1}{a}}{x-a}\] \[\Large \frac{ax(\frac{1}{x}) - ax(\frac{1}{a})}{ax(x)-ax(a)}\] \[\Large \frac{a-x}{ax^2-a^2x}\] \[\Large \frac{a-x}{-ax(a-x)}\] \[\Large \frac{\cancel{a-x}}{-ax\cancel{(x-a)}}\] \[\Large -\frac{1}{ax}\] So \[\Large \frac{f(x) - f(a)}{x-a}=-\frac{1}{ax}\] when \[\Large f(x)=\frac{1}{x}\]

jimthompson5910 (jim_thompson5910):

\[\Large f(a) = \frac{1}{a}\] because all that we did was replace 'x' with 'a'

OpenStudy (anonymous):

how do you come across multiplying the numerator and denominator by ax/how do you determine that's the common denominator?

jimthompson5910 (jim_thompson5910):

the LCD of 1/x and 1/a is ax, multiplying EVERY term by this clears out the inner fractions

jimthompson5910 (jim_thompson5910):

the LCD is the LCM of the denominators to find the LCM of a and x, just multiply them and then divide by the GCF. Since the GCF is 1, we get (ax)/1 = ax